OpenStudy (anonymous):

x-14 squarerootx +48=0

OpenStudy (anonymous):

u=sqrt(x) so u^2 -14u +48 =0

OpenStudy (anonymous):

using factorization method u^2-12u-4u+48=0 (u-12)(u-4)=0 u=12 or u=4

OpenStudy (anonymous):

(u-8)(u-6) =0

OpenStudy (anonymous):

the answer has to be solutions

OpenStudy (anonymous):

oh u right man

OpenStudy (anonymous):

u=6 , u=8 sqrt(x) = 6 --> x = 36 sqrt(x) = 8 --> x= 64

OpenStudy (anonymous):

thanks how did you get the answer please tell me how

OpenStudy (anonymous):

once u find the values of u, by backward substitution u find x

OpenStudy (amistre64):

you simply see how it could fit into a "normal" quadratic form. if u = sqrt(x) then u^2 = x right? So we solve for: u^2 - 14u +48 = 0 instead; and when we find the values for "u" we can then determine that value for x; since u = sqrt(x) when "u" = 6 we get: 6 = sqrt(x); or 6^2 = x, so naturally, x = 36 for this solution

OpenStudy (amistre64):

x -14 sqrt(x) +48 = 0 | | | | u^2 - 14 u +48 = 0